If population mean and standard deviation is 100 and 10 respectively. And sample mean is 115 and n = 30 then what is z score?


1
Expert's answer
2021-06-01T17:48:38-0400

Z score formula for the mean of n observation


"\\z=\frac{\bar x-\mu}{(\frac{\sigma}{\sqrt{n}})}"


"\mu =100\\
\sigma=10\\

\bar x=115\space \\ n=30"


"\\z=\frac{115-100}{(\frac{10}{\sqrt{30}})}\\
\\z=\frac{15}{1.82574186}\\
z=8.21584061"